I am not sure how to solve this equation.
(3/4-2/3)+ (1/2+1/3)=

I also need help solving this equation.
1/2+(2/3 divide by 3/4)-(4/5x5/6)=

Answer by vertciel(183)

Hello there,
The first question:
(3/4 - 2/3) + (1/2 + 1/3)
= 3/4 - 2/3 + 1/2 + 1/3 <-- Remove bracket
= -1/3 + 1/2 + 3/4 <-- Simplify what you can
= (-4 + 6 + 9)/12 <-- Find the Lowest Common Multiple LCM and set the denominator to it.
= 11/12

With regards to your second question, I don't understand your third bracket. Could you please write it more clearly?